### v2

# Movie2Story: A framework for understanding videos and telling stories in the form of novel text

Abstract
In recent years, large-scale models have achieved significant advancements, accompanied by the emergence of numerous high-quality benchmarks for evaluating various aspects of their comprehension abilities. However, most existing benchmarks primarily focus on spatial understanding in static image tasks. While some benchmarks extend evaluations to temporal tasks, they fall short in assessing text generation under complex contexts involving long videos and rich auxiliary information. To address this limitation, we propose a novel benchmark: the Multi-modal Story Generation Benchmark (MSBench), designed to evaluate text generation capabilities in scenarios enriched with auxiliary information. Our work introduces an innovative automatic dataset generation method to ensure the availability of accurate auxiliary information. On one hand, we leverage existing datasets and apply automated processes to generate new evaluation datasets, significantly reducing manual efforts. On the other hand, we refine auxiliary data through systematic filtering and utilize state-of-the-art models to ensure the fairness and accuracy of the ground-truth datasets. Our experiments reveal that current Multi-modal Large Language Models (MLLMs) perform suboptimally under the proposed evaluation metrics, highlighting significant gaps in their capabilities. To address these challenges, we propose a novel model architecture and methodology to better handle the overall process, demonstrating improvements on our benchmark.
Movie2Story: A framework for understanding videos and telling stories in the form of novel text
Kangning Li kangningl25@ucla.edu                        Zheyang Jia u3621320@connect.hku.hk                        Anyu Ying anyuy@andrew.cmu.edu
1 Introduction
In recent years, many multi-modal large language models (MLLMs)Alayrac et al. (2022); Zhu et al. (2023); Li et al. (2024a); Huang et al. (2023); Li et al. (2020) have effectively combined video encodersTouvron et al. (2023); Devlin et al. (2019); Dosovitskiy et al. (2021) and audio encodersRadford et al. (2022); Chen et al. (2022b, a) to achieve strong performance in vision-based description generation and audio-based tasks. However, as MLLMs evolve, a critical challenge arises: how can we rigorously evaluate their comprehension and generation capabilities? Addressing this question is vital for assessing current models and guiding future developments.
Existing benchmarks for MLLMs primarily adopt a question-answering (QA) formatYu et al. (2024); Xu et al. (2023); Xie et al. (2024); Fu et al. (2024a); Xu et al. (2016); Liu et al. (2024); Cheng et al. (2024), focusing on static image understanding. While benchmarks such as MVBenchLi et al. (2024b) extend evaluation to temporal tasks, they remain inadequate for assessing long-video comprehension and fail to incorporate rich auxiliary information such as audio. Although some studies, such as AV-SUPERBTseng et al. (2024), attempt to integrate audio information with visual information, they primarily focus on the evaluation of audio. Moreover, current benchmarks typically produce objective, template-like outputs, lacking the stylistic complexity of narrative storytelling. Furthermore, many benchmarks rely heavily on manual annotations, which are resource-intensive and time-consuming.
To overcome these limitations, we introduce a novel benchmark, Multi-modal Story Generation Benchmark (MSBench), which emphasizes the integration of diverse auxiliary information (e.g., audio features) to generate long-video descriptions in a narrative and information-rich style.
Our approach features an innovative data transformation pipeline tailored to enhance existing video-text datasetsLi et al. (2020). Recognizing the brevity and limited scope of current datasets, we augment them with detailed audio and visual information. For instance, simple descriptions like "a car is driving" are enriched into nuanced narratives such as "a man sings: ‘…’, while a woman drives the car…". Auxiliary information, including audio-derived attributes such as ASR (Automatic Speech Recognition)Kheddar et al. (2024) outputs, emotions, and sound events, is extracted using powerful open-source audio models. For visual descriptions, we leverage state-of-the-art image understanding modelsBai et al. (2023); Chen et al. (2024) to generate detailed frame-level descriptions, ensuring narrative consistency through temporal ordering.
Our benchmark offers two significant advantages. First, it drastically reduces the reliance on manual annotation by utilizing automated processes and open-source tools. Second, it introduces five novel evaluation metrics tailored to assess text quality in enriched information scenarios. These metrics span dimensions from environmental and psychological descriptions to name-entity specificity, providing a comprehensive assessment of narrative text quality in open-world, long-duration contexts.
Finally, we evaluate several state-of-the-art MLLMsLi et al. (2024a); Lin et al. (2024); Wang et al. (2022, 2024) on the MSBench benchmark, revealing substantial gaps in their performance. Models like VideoChat2Li et al. (2024b) struggle with incorporating audio features, handling long-video contexts, and generating high-quality narrative outputs. To address these challenges, we propose a robust baseline model, M2S-LLM, specifically designed to align with our benchmark’s requirements and enhance text generation quality. Experimental results demonstrate that M2S-LLM surpasses VideoChat2Li et al. (2024b) by nearly 15% across key metrics.
All models, datasets, and evaluation frameworks are publicly available to facilitate future research and advancements in the field.
2 Related Works
2.1 MLLM Introduction
The evolution of Large Language Models (LLMs) has accelerated research into Multi-modal Large Language Models (MLLMs)Lyu et al. (2023); Lee et al. (2024); Fu et al. (2024b), aiming to integrate diverse modalities like text, vision, and audio. Early breakthroughs, such as FlamingoAlayrac et al. (2022) and PaLM-EDriess et al. (2023), showcased strong performance in multi-modal tasks by combining visual and textual modalities. Subsequent open-sourced efforts, including LLaVALiu et al. (2023) and MiniGPT-4Zhu et al. (2023), have expanded the scope of multi-modal instruction tuning, while VideoChatLi et al. (2024a) and VideoChatGPTMaaz et al. (2024) extended these ideas to dynamic video tasks by utilizing ChatGPT-generated annotations.
Building on these foundations, several recent works have introduced innovative approaches to tackle the complexities of video comprehension. The MMADYe et al. (2024b) model integrates video, audio, and textual information to generate concise, real-time descriptions. By leveraging pre-trained encoders and advanced fusion techniques, MMADYe et al. (2024b) enhances narrative richness and accessibility. Similarly, the framework for Distilling Vision-Language Models on Millions of VideosZhao et al. (2024) adapts image-language models for video-language tasks, generating large-scale pseudo-labels with multi-granularity features, offering enhanced semantic and contextual understanding.
The Video-LLaMA2Cheng et al. (2024) model aligns audio, video, and textual data into a unified vector space using the pre-trained ImageBindGirdhar et al. (2023) module. Instead of training an audio-text dataset, it utilizes a video-text encoder, indirectly converting audio into text. Meanwhile, Video StorytellingLi et al. (2020) focuses on generating textual summaries for events by selecting key frames with a reinforcement learning-based Narrator model and employing contextual embeddings through Residual Bidirectional RNN (ResBRNN)Li et al. (2020), with Transformers enhancing long-range dependency modeling for richer, more cohesive stories.
Our work builds upon these advancements by addressing the limitations in existing MLLMs, particularly in handling long-duration videos, integrating auxiliary information, and generating stylistically rich narrative outputs. By combining innovative dataset enhancement techniques and robust evaluation metrics, our framework sets a new standard for multi-modal understanding and narrative-driven generation.
2.2 Benchmark Introduction
Traditional Vision-Language (VL) benchmarksGoyal et al. (2017); Kay et al. (2017); Xu et al. (2016, 2017); Xiao et al. (2021) have predominantly focused on QA-style evaluations, targeting tasks like multi-modal retrieval and vision QA. More recent benchmarks have broadened the scope to assess integrated capabilities. For example, OwlEvalYe et al. (2024a) and SEED-BenchLi et al. (2023) introduced evaluation metrics that transcend simple model ranking, focusing on comprehensive multi-modal skills. In the video domain, benchmarks like Perception TestPătrăucean et al. (2023) examine multi-modal video perception and reasoning, while FunQAXie et al. (2024) pushes reasoning boundaries using humorous or counter-intuitive content.
MVBenchLi et al. (2024b) stands out by defining over 20 tasks to evaluate MLLMs’ performance on diverse scenarios, especially temporal reasoning. The study on Synchronized Video StorytellingYang et al. (2024) demonstrates an innovative approach by utilizing additional ad keywords for generating and assessing the quality of storytelling and advertisement content, providing considerable inspiration for the field. Similarly, MMADYe et al. (2024b) highlights the evaluation of critical but non-automatically measurable elements, such as character names, enriching benchmark design. These advancements inform our framework, which aims to address gaps in evaluating narrative-driven, multi-modal tasks.
3 MSBench
In this section, we delve into the details of our MSBench. Initially, we design the multi-modal story generation tasks as illustrated in Figure 1. Subsequently, we automatically generate caption-story pairs for evaluation, as depicted in stage 2 and outputs in Figure 1.
3.1 Story Task Definition
To craft the story tasks for MSBench, we employ a text-to-text methodology, which involves integrating scattered text information into long texts in the form of novels.
As previously highlighted in the introduction, the majority of current MLLM benchmarks are centered around direct transformations from video to text or audio to text, often resulting in brief summary captions. Driven by this observation, we propose leveraging these existing task definitions as a foundation to systematically develop multi-modal story generation tasks.
We begin by summarizing the main tasks of video understanding and audio understanding text summaries from previous benchmarks. Then we enrich these tasks to longer stories, creating story tasks that can not be effectively solved with a single model (video or audio) and require comprehensive understanding of video and audio information. Finally, we define 5 tasks as follows.
video captions. A comprehensive understanding for video information, generated by VQA model like videochat2Li et al. (2024b).
Audio ASR. Text version of the character’s speech in the audio, exactly what they said in the audio.
Audio Feature. Emotion and speech speed summary for characters. And speaker recognition.
Video Audio Alignment. Matching characters in movie.
Timestamp matching and Story line Processing. Timestamp alignment for video and audio information and generation of storyline.
3.2 Automatic caption-story pairs Generation
M2S maintains a Foundation Models Pool to store various video and audio foundation models, which own the capability to detect and annotate different features in videos. After having sufficient features, we use LLM to integrate the features in chronological order and generate a lengthy novel level story text.
M2S maintains a robust Foundation Models Pool. These models are equipped with the advanced capability to detect and annotate a myriad of features within videos. Once a sufficient array of features has been collected, we leverage a Large Language Model (LLM) to systematically integrate these features in chronological order (sorted by "timestamp" or segments as shown in Figure 1 video/audio captions respectively), thereby generating a comprehensive, novel-level story text that is both lengthy and rich in detail.
3.2.1 Data Pre-processing
(1) Video splitting: The video may have a long length, and a short 20s video is enough to generate a long story text. Therefore, dividing a long video into 20s has the following advantages: 1. Increasing the amount of data 2. Unifying the data size 3. Reducing memory usage. (2) FFmpeg: A Video Frame Extraction Tool, used for extracting key frames from videos.FFmpeg Developers (2023)
3.2.2 Video-Audio Feature Information Generation..
In our pipeline, feature extraction can be separate to several parts, here we list these models for extracting desired features and generating corresponding texts for our story tasks proposed in previous subsection.
video captions. Use Pre-trained Models(video-language model), including MMADYe et al. (2024b) and VideoChat2Li et al. (2024b) for generating detailed descriptions and dialogue content.
Audio ASR. Use speech Extraction Tool, eg. WhisperOpenAI (2022) model for speech-to-text conversion, noting it generates text without capturing emotions.
Audio Feature. Use Emotion Analysis Module Using BERT-like encoder networks for audio emotion analysis to extract emotional information and transform in text version, eg Emotion2Vec.et al. (2023c)
Video Audio Alignment. Character Relationship Network. Use CNNs like YOLOv8Yaseen (2024) for character recognition and audio processing models (eg.pyannote)et al. (2019) to split different characters. Matching them by anchors predefined by human.
Timestamp matching and Story line Processing. Using audio timestamp splitting and video key frame resampling techniques to link information across different segments.
3.2.3 Story Generation:
Using Large Language Models, such as GPT-4oet al. (2024b), used to expand and refine generated descriptions into a complete novel. Here we use llama3-3Bet al. (2024a), mistral7Bet al. (2023a), Qwen7Bet al. (2023b) Collect all information and time sequence of video-audio text descriptions and generate a coherent story by well designed instructions and prompts.
3.3 Prompt Design for Evaluation
To emphasize the story line coherence of MLLMs, we craft a detailed system prompt for evaluation (see the top right of Figure 1). This prompt encourages MLLMs to carefully scrutinize video and audio content to generate stories, by paying attention to factors such as timestamps.
3.4 Additional caption-stroy pair for fine-tuning LLMs
We use GPTZhu et al. (2023) to generate virtual data in the corresponding format to enhance data quality. For example, as shown in Figure 1, first provide the format and example of the corresponding data for GPT. Then let GPT generate data in the corresponding format as virtual data and add it to the dataset for LLMs to make fine adjustments. Because at this stage, we only need to adjust the long text generated in the form of a novel, focusing mainly on whether the generated story contains the information in the caption and temporal alignment, rather than whether the information in the caption is in the original video and audio. Therefore, the data generated in this way is reasonable.
4 M2S pipeline
Figure 1: Workflow template
After building our MSBench, we evaluate a number of popular video caption MLLMs in the first three rows in Table 2. Surprisingly, the existing MLLMs are far from satisfactory in temporal understanding. To fill the gap, we develop a robust MLLM-LLM pipeline, which is dubbed as M2S(movie to story). The basic process is the same as the way MSBench generates data. The basic process is the same as the way MSBench generates data. Here are some details of Video Audio Feature Information Generation and LLMs. The text integrated through LLM generally has a lower repetition rate and higher information score (column 2 and 3 in Table 2.)
4.1 Video Feature Information Generation
Our video processing pipeline first segments the video length, typically 20 seconds, to facilitate efficient processing and analysis. Then use open-source VLM and other models to extract video/image text descriptions from video files. The data is stored in JSON format for downstream text processing, which includes not only video content but also timestamps. Therefore, subsequent LLM can process it to obtain highly consistent text content.
Video captions Our approach employs a segmented video processing strategy, which effectively addresses the challenges of handling long videos while maintaining the continuity of the narrative. Within each video segment, we focus on extracting storyline information and background context as the key components of the video segment. In terms of model selection, we conducted a comprehensive evaluation of various open-source models, including videollavaLin et al. (2024) and InternvideoWang et al. (2022) . Following a comparative analysis, we finally adopted for the open-source model Videochat2Li et al. (2024b) as our baseline model. By carefully designing the prompt format for VQA models, we ensured that the model could accurately extract the necessary key information, laying the foundation for in-depth analysis of the video content.
Image Character Recognition Our goal is to match real-person names with video characters by recognizing and matching individuals across frames in long videos, despite challenges like varying appearances and multiple viewpoints. We use FacenetSchroff et al. (2015) for robust facial recognition and a text feature-based method with VITDosovitskiy et al. (2021) and LLM for short-segment matching. For global matching, we establish a dictionary of key characters with face and text features as anchors. YOLOv8Yaseen (2024) is used for pre-segmenting person bounding boxes, and we filter main characters by comparing anchor-candidate similarity. This approach combines vector and text features to effectively match individuals across scenes in long videos.
4.2 Audio Feature Information Generation
Our audio processing pipeline begins with the extraction of audio from video files using Python’s audio libraries. The extracted audio is then segmented into smaller, manageable chunks, facilitating efficient processing and analysis. The data is stored in JSON format for downstream text processing, which includes not only audio features and content but also timestamps. Therefore, it can be processed by subsequent LLMs to obtain highly consistent text content.
Speaker Embedding for speaker recognition The core of our speaker recognition system lies in the Speaker Embedding module, powered by pyannote.audioet al. (2019). This module leverages deep learning models to extract unique and discriminative features from each speaker’s voice. These speaker embeddings serve as compact representations that capture the essential aspects of a speaker’s identity, facilitating accurate recognition and matching.
Emotion Analysis Emotion2Vecet al. (2023c) is an innovative deep learning based emotion recognition tool that aims to encapsulate the essence of emotional vocabulary in a rich high-dimensional vector space. This method not only captures the subtle relationships between words, but also assigns them the emotional burden of human expression. Emotion2Vec demonstrated the power of deep learning in the field of emotion recognition, providing emotional storytelling through vector language. The model in this article is derived from the Emotion2Vec related model trained by FunASR.
Word Speed This part is relatively simple, just calculate the number of words spoken by the speaker in a unit of time to obtain the speaking speed, using text descriptions such as "fast" and "slow".
4.3 LLMs selection and LoRA Finetuning
Different LLMs have varying abilities to integrate information and respond differently to instructions. Moreover, as a proprietary task, generic LLM may not necessarily have good results, so techniques such as LoRA may be needed to fine tune it. After extracting video and audio information, we can fine tune LoRAet al. (2021) for different LLMs to complete the final stage of generation and obtain stories.
Note that the data here can be real video audio caption story pairs generated through our pipeline. As introduced in the previous section, GPT can also be used to generate virtual data in the corresponding format to enhance data quality.
5 Experiment
5.1 Quantitative Analysis
Traditional NLP metrics such as BLEUPapineni et al. (2002), ROUGELin (2004) evaluate generated text by comparing it to reference ground truth descriptions. While effective for assessing surface-level language quality, these metrics fail to capture the richness of knowledge beyond the visual domain, such as audio-related information, character details, or contextual nuances.
For instance, while a visual scene may depict a list of ingredients, a narration could highlight only a subset of those ingredients or incorporate additional elements like speaker emotions or environmental sounds, which are critical for generating a more comprehensive story. Since traditional metrics rely solely on visual references, they do not account for these extra layers of information.
To address this gap, we propose a set of reference-free metrics: Language Fluency, Key-knowledge Relevance. These metrics evaluate fluency, knowledge integration (including both visual and audio elements), and narrative alignment, providing a more accurate assessment of the generated stories and better reflecting the advantages of our multi-modal framework.
Language Fluency: Fluency is used to evaluate the coherence and sentence repetition rate of generated explanatory texts. To assess language fluency, we utilize a keyword-triplet method that evaluates fluency at three levels of granularity: (1) within a single sentence, (2) between different sentences, and (3) the repetition of key-knowledge triplets relative to the overall story. Repetition of triplets within the story is considered a sign of unnatural phrasing and redundancy, which negatively impacts fluency. A higher fluency score is assigned when the text flows smoothly, with minimal repetition and a coherent structure. We have made modifications to the formulas in the original paperYang et al. (2024).Fluency and repetition are opposite. In the chart, we use repetition to represent the fluency of language, with lower repetition indicating smoother language. Then repetition can be represented by Intra-story Repetition (ISR): $\text{ISR}(S)=\frac{1}{n(n-1)}\sum_{i\neq j}\frac{\left|W(s_{i})\cap W(s_{j})\right|}{\left|W(s_{i})\right|}$ where $n$ is the number of explanatory texts, and $W(s_{i})$ represents the vocabulary set of explanatory text $s_{i}$. The formula calculates the repetition rate of vocabulary within the explanatory texts, which should be minimized to ensure diversity and coherence in the text.
Key-knowledge Relevance: For evaluating the relevance of key knowledge in the generated story, we define a knowledge base that includes both visual and audio information. In contrast to traditional models that focus solely on visual knowledge, our approach incorporates additional audio-related knowledge, such as ASR (Automatic Speech Recognition), environmental sound events, and other audio features. This expanded knowledge base allows us to assess the relevance and diversity of knowledge within the story. Specifically, we use two evaluation criteria: - Information Similarity (InfoSim)Yang et al. (2024): Measures the alignment between the knowledge points in the story and the knowledge repository. A higher similarity score indicates that the generated story effectively incorporates relevant knowledge. - Information Diversity (InfoDiverse)Yang et al. (2024): Evaluates the breadth of the knowledge used in the story. A higher diversity score indicates that the story incorporates a wide range of knowledge points, avoiding over-reliance on a small subset of information. The formula here is quite different from the original text. We have changed the normalization for time to the normalization for the number of keywords. That is to say, InfoDiverse calculates the proportion of keywords in the generated data that match the keywords in the reference data, ranging from 0 to 1.
InfoSim $=\frac{1}{2|s_{i}|}\sum_{s_{i}}\left(\max_{k\in K}f_{k}^{T}f_{s_{i}}\right.$
$\quad\left.+\frac{1}{|W(s_{i})|}\sum_{w\in W(s_{i})}\max_{k\in K}f_{k}^{T}f_{w}\right)$ (1)
InfoDiverse $=\frac{1}{|K|}\left|\bigcup_{s_{i}}\{k_{t}\in K\mid\max_{w\in W(s_{i})\cup S_{i}}\right.$
$\quad\left.f_{k_{t}}^{T}f_{w}>0.9\}\right|$ (2)
where $W(s_{i})$ represents all words in sentence $s_{i}$. $f_{s_{i}}$, $f_{k}$, and $f_{w}$ refer to the normalized embeddings of sentence $s_{i}$, knowledge point $k$, and segmented word $w$, respectively.
A well-rounded story should demonstrate both high information similarity and diversity, ensuring that relevant visual and audio knowledge are integrated in a meaningful way.
These evaluation metrics provide a comprehensive framework for assessing the quality of generated story descriptions, ensuring that the generated narratives meet high standards in terms of fluency, knowledge relevance, and visual coherence.
5.2 Qualitative Analysis
5.2.1 GPT Assistant Metrics
In this section, we introduce a set of key evaluation metrics designed to assess the quality of generated story descriptions using GPT and prompt engineering. By defining these metrics, we aim to leverage GPT as a rigorous evaluator of novel-like text, assigning scores based on specific story elements. These metrics are particularly useful for assessing narrative quality in terms of various aspects such as environmental details, emotional expression, language proficiency, character portrayal, and storyline coherence.
The following metrics are defined:
Environment Description Score: This score evaluates how well the generated text sets the scene and conveys the environment. Key factors include the richness of the setting, sensory detail, and atmosphere. A high score indicates a vivid and immersive portrayal of the environment.
Emotional Description Score: This metric assesses the depth and authenticity of emotional expression in the story. It measures how well the text captures and conveys the emotional states of characters, as well as the emotional tone of the narrative.
Language Description Score: This score focuses on the clarity, fluency, and grammatical quality of the text. It evaluates how well the story is written, including sentence structure, vocabulary choice, and coherence. A higher score reflects a more polished and readable narrative.
Character Description Score: This metric evaluates how effectively characters are described in terms of appearance, personality, and behavior. It gauges whether the character descriptions are detailed, nuanced, and consistent throughout the narrative.
Structural and Rationality Coherence Index (SRCI) Score: This metric quantifies the narrative coherence and the logical progression of the storyline. It gauges the effectiveness of plot construction, the lucidity of the story’s trajectory, and the seamless incorporation of essential storytelling components such as conflict, resolution, and character progression. The SRCI reflects the story’s ability to maintain a consistent and engaging structure that resonates with the audience’s expectations of narrative rationality.”
Storyline Accuracy with Canonical Outline Reference (SACOR) Score: This score measures the accuracy of the generated story against a predefined canonical storyline. It assesses how closely the generated narrative aligns with the key plot points, character behaviors, and thematic elements of the original story outline. The SACOR score is indicative of the model’s fidelity to the source material and its ability to reproduce the intended narrative arc with precision.”
For a more comprehensive evaluation, these scores are tabulated and compared across different story generations. The results can be found in Table 4, which summarizes the performance of our framework on each metric.
In Table 4, the scores represent the average evaluation across multiple generated stories, providing a comprehensive view of the performance of our framework compared to baseline methods.
5.2.2 Manual Metrics
Table 4 shows the results of applying our M2S framework for generating detailed story descriptions from several movies. The comparison between our framework and baseline methods demonstrates the clear advantages of our approach.
More specifically, our framework excels in producing more comprehensive and accurate story descriptions. By leveraging the Audio module, we incorporate key information such as ASR (Automatic Speech Recognition) output, speaker identity, and speaking rate to further enhance the generated descriptions. Additionally, the combination of the Actor-Matching module, Storyline-Segment module, Audio module, and Vision module enables our framework to generate rich, long-form narrative descriptions that go beyond traditional subtitles, offering a novel way of understanding movies. This integrated approach results in more detailed and accurate descriptions, which could serve as an independent means of movie comprehension, rather than merely being a supplementary subtitle track.
6 Conclusion
In this paper, we introduced MSBench, a comprehensive benchmark designed to evaluate the multi-modal story generation capabilities of MLLMs. We also proposed a robust MLLM pipeline, M2S, which outperforms leading models on the MSBench benchmark. Our extensive analyses provide valuable insights into the design of MLLMs for multi-modal story generation, particularly in scenarios enriched with additional information. Despite these advancements, our current approach has some limitations. We aim to address these in future work to enhance the evaluation framework and further refine the performance of MLLMs in complex, information-rich environments.
Appendix A Appendix Tables
Model ISR Info_sim Info_diverse SACOR Rouge-1 Rouge-2 Rouge_L BLEU-4
Internvideo2 15.1
VideoLlava 42.7
Videochat2 8.1 38.0 55.0
Internvideo2 + GPT4o 4.0 26.5 18.6 59.4 39.6 10.2 18.6 2.7
VideoLlava + GPT4o 5.6 25.0 20.7 61.1 40.9 10.5 19.0 3.2
Videochat2 + GPT4o 4.5 27.2 18.3 61.2 40.1 10.1 18.8 2.9
Internvideo2 + A-GPT4o 5.6 25.5 17.6 67.0
VideoLlava + A-GPT4o 6.0 25.2 21.9 65.3
Videochat2 + A-GPT4o 12.4 27.3 53.6 52.8
Table 1: Performance of various models with GPT4o as LLM on Quantitative metrics
Model Environment Character Emotion Language SRCI Overall GPT Human
Internvideo2 2.494 1.525 0.608 1.825 1.421 1.575
VideoLlava 2.170 1.155 0.213 1.539 1.118 1.239
Videochat2 3.086 1.878 0.921 2.243 1.809 1.987
Internvideo2 + GPT4o 3.766 2.675 1.792 3.555 2.717 2.901
VideoLlava + GPT4o 3.592 2.586 1.667 3.438 2.594 2.775
Videochat2 + GPT4o 3.873 2.651 1.738 3.550 2.727 2.908
Internvideo2 + A-GPT4o 3.581 2.574 1.659 3.433 2.585 2.767
VideoLlava + A-GPT4o 3.427 2.491 1.607 3.195 2.485 2.641
Videochat2 + A-GPT4o 3.492 2.568 1.673 3.090 2.625 2.689
Table 2: Performance of various models with GPT4o as LLM on Qualitative metrics
Model ISR Info_sim Info_diverse SACOR Rouge-1 Rouge-2 Rouge_L BLEU-4
GPT4o 12.4 27.3 53.6 52.8 - - - -
GPT3.5 13.5 26.3 25.5 54.0 38.2 8.6 17.6 1.9
Doubao 5.4 27.7 16.0 56.8 32.6 7.0 15.7 0.8
QWen 3.4 33.2 46.3 60.1 46.2 17.8 22.3 5.8
QWen-7B-instruct 9.8 35.3 50.8 58.3 21.4 34.8 16.5 27.3
mistral-instruct-7B-v1.0 8.1 8.3 7.5 3.5 4.7 1.8 6.3
llama3-3B 22.1 - 7.9 16.6 8.6 21. 24.6 8.2
lora_QWen-7B-instruct 7.6 33.1 50.8 58.2 52.4 21.7 25.1 9.0
lora_mistral-instruct-7B-v1.0 7.3 33.0 50.1 58.0 52.2 21.7 25.0 8.9
lora_llama3-3B 8.5 33.0 52.4 58.8 52.2 21.2 24.6 8.7
Table 3: Performance of various LLMs on Quantitative metrics
Model Environment Character Emotion Language SRCI Overall GPT Human
GPT4o 3.492 2.558 1.673 3.090 2.625 2.689 -
GPT3.5 2.831 1.809 1.077 2.299 1.815 1.966 -
Doubao 2.998 1.906 1.025 2.233 1.878 2.008 -
QWen 3.491 2.356 1.567 2.994 2.376 2.557 -
QWen-7B-instruct 2.826 1.884 1.125 2.224 1.805 1.973 -
mistral-instruct-7B-v1.0 0.360 0.229 0.124 0.297 0.213 0.245 -
llama3-3B 0.294 0.187 0.084 0.261 0.134 0.192 -
lora_QWen-7B-instruct 3.606 2.468 1.611 3.170 2.504 2.672 -
lora_mistral-instruct-7B-v1.0 3.623 2.494 1.625 3.159 2.527 2.686 -
lora_llama3-3B 3.560 2.431 1.577 3.106 2.457 2.626 -
Table 4: Performance of various LLMs on Qualitative metrics